OpenStudy (anonymous):

Explain the difference between using the tangent ratio to solve for a missing angle in a right triangle versus using the cotangent ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (anonymous):

@nirmalnema PLEASE HELP ME

OpenStudy (anonymous):

@ikram002p THANKS SO MUCH

OpenStudy (ikram002p):

let us have this right angle 

OpenStudy (anonymous):

ok

OpenStudy (ikram002p):

so if u need to find angle B , and you know measure of sides a,b how would u find it ?

OpenStudy (ikram002p):

tan B =c/b or cot B = b/c

OpenStudy (anonymous):

cot?

OpenStudy (anonymous):

no! tan

OpenStudy (ikram002p):

whe u start with finding tangent ratio , tan B =c/b B =tan^-1 (c/b) how ever if u wanna use cotangent ratio cot B = b/c B =cot^-1 b/c or note that cot B= 1/ tan B , so cot B=1/tan B = 1/b/c =c/b thus b=tan^-1 (c/b) @ganeshie8 is there anything we should explain ?

OpenStudy (anonymous):

So how could I explain that in words? I still don't understand a lot of it.

ganeshie8 (ganeshie8):

There is not much difference because both ratios require you to know the same sides (opposite and adjacent) : tan(B) = opposite/adjacent cot(B) = adjacent/opposite So the only difference is that one ratio is the reciprocal of the other ratio...

OpenStudy (anonymous):

Thank you! :)

OpenStudy (ikram002p):

oh ic ,but i would say using tan is much better since maybe u use calculator , so u would reduce one step ?!

ganeshie8 (ganeshie8):

yes you should add that step also... because calculators come with tan ratio only

OpenStudy (ikram002p):

yeah i guess , idk about new calculators though